Angelo, Brandon, and Carl work in the same office. Angelo's age is 4 years more than twice Carl's age. Brandon is 5 years younger than Carl. The average of the 3 ages is 41. Find the age of each of the men.
Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(48546)

Angelo, Brandon, and Carl work in the same office. Angelo's age is 4 years more than twice Carl's age. Brandon is 5 years younger than Carl. The average of the 3 ages is 41. Find the age of each of the men.
--------------------
Let Angelo be a; Brandon is b ; Carl is c
EQUATION:
a = 2c+4
b = c-5
(1/3)(a+b+c) = 41
-----------------
Rearrange:
a + 0 - 2c = 4
0 + b -c --5
a + b + c = 123
-------------------
Using the Matrix function of a TI calculator I get:
a= 66
b= 26
c= 31
==========
Cheers,
Stan H.

Answer by ankor@dixie-net.com(12694)